What is the surface area of a sphere with volume 10.67 pie m cubed. Give your answer in terms of pie please and thank you
Answer by mananth(16946)

4/3*pi^r^3=10.67
3*10.67 / 4pi = r^3
25.14 =r^3
r=2.93
surface area = 4*pi*r^2
= 4*2.93^2*pi
=34.33pi